OpenStudy (anonymous):

Brett has determined a function f(x) that shows the exponential growth of the number of shoes Larae owns each year. Explain how the f^−1(x) can be found and what f^−1(132) means.

OpenStudy (anonymous):

f(x) . f−1(x) is the inverse function of f(x), it can be found by replacing the x and y values like if f(x) = x+2 y = x+ 2 inverse is x = y + 2 x - 2 = y x-2 = f^-1(x)..

OpenStudy (anonymous):

but what values am i supposed to be replacing?